Which of the following proportions is TRUE???
a) 1:4 = 2:5
b) 6:9 = 10:12
c) 5:11 = 15:30
d) 2:7 = 4:14

Answer by Alan3354(69443)

Which of the following proportions is TRUE???
a) 1:4 = 2:5
1/4 = 2/5 ?
Cross multiply
1*5 = 4*2 ?
Not true.
Do the others the same way.
------------------------------
I don't usually answer questions with more than one question mark (?).
What do you think ??? accomplishes?
And, all CAPS is something I avoid.
Math is no place for hysteria.
-----------------------------
b) 6:9 = 10:12
c) 5:11 = 15:30
d) 2:7 = 4:14